OpenStudy (anonymous):

1- y*lnx*lny*dx+x*dy=0

OpenStudy (amistre64):

are these exact equations?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then how much do you know about finding a solution

OpenStudy (anonymous):

this is equations .. how i can solution?

OpenStudy (amistre64):

the general process is to integrate one side, add on a function, and derive to the other side to compare with

OpenStudy (amistre64):

given: M(x,y) dx + N(x,y) dy = 0 integrate one side int: M(x,y) dx add a function: int: M(x,y) dx + f(y) and take the derivaitve d/dy [int: M(x,y) dx + f(y)] = N(x,y) dy

OpenStudy (amistre64):

y*lnx*lny*dx+x*dy = 0 since xdy is simple enough .... integrate to get xy add a function of x to it: xy+ f(x) and take the derivative, equate it to the other side [xy+ f(x)]' = y + xy' + f'(x) = y*lnx*lny*dx maybe the other direction?

OpenStudy (amistre64):

i got no good way past that .... see if you can glean something from the wolfs http://www.wolframalpha.com/input/?i=y*lnx*lny*dx%2Bx*dy+%3D+0

OpenStudy (anonymous):

Thank you very much

OpenStudy (amistre64):

yep, good luck :)